# A Study on Asynchronous Vote-based Blockchains

### v1

@@ 5.3.2 Sending inconsistent blocks
6 Optimal Two-Round Voting Process
HotStuff-2 requires three continuous block intervals to reach a consensus. During these intervals, all honest nodes are aligned, and messages are successfully exchanged before each interval concludes. Similarly, our protocol also necessitates three block intervals: one for proposing the block and two for voting. In the initial block interval (consensus height $k^{\prime}$), block $B_{k^{\prime}}$ is proposed and received by all before the interval ends. In the subsequent interval, all honest nodes cast their votes for $B_{k^{\prime}}$, and these votes are received by all before the interval’s close. By the third interval (vote height $k^{\prime}+2$), all honest nodes can verify through their vote counts that $B_{k^{\prime}}$ has garnered enough support during vote height $k^{\prime}+1$, thereby satisfying Eq. 2 and achieving consensus. Thus, our protocol aligns with the optimal voting rounds of HotStuff-2 and efficiently supports $O(N)$ view changes.
@@ 7 Implementation
We did not experiment with DAG-based protocols. They exhibited similar complexity to ACS-based protocols.
@@
At the start of each 5 seconds, 1000 identical transactions (each sized 256 bytes) in total appeared in each node’s local database. This was regardless of the progress of the blockchains. The leader node grouped the first 1000 transactions not written to the blockchain to form a block. For nodes in Dumbo, they broadcasted the uncommitted transactions, which together sized $2.05\times 10^{6}$ bits.
@@ 7.1.1 Methodology
Regarding voting, adversarial nodes in our algorithm were constrained to make compliant votes. If they voted, they expedited consensus. We, therefore, assumed that adversarial nodes in our protocol and Hotstuff would attempt to create forks when selected as a leader, and otherwise kept silent, akin to the behavior of silent adversaries in Dumbo.
@@ 7.1.2 Results
Our protocol used a similar average block acceptance time compared to HotStuff-2, while HotStuff-2 required smaller data; both approaches largely outperformed Dumbo2 and Dumbo-NG due to reduced complexity, as expected. Tbl. 1 reported the experiment results.
@@ 7.2.2 Results
As expected, our protocol largely outperformed Dumbo2 and Dumbo-NG due to reduced complexity. Tbl. 2 presented the experimental results. The transaction inclusion rate for our protocol was similar in both good and bad-case latency because the only factors that affected it in our experiment were the rate of adversarial leaders and the pre-set block interval. The average data usage per node per instance increased because rounds of voting occurred while awaiting the votes from the slowest honest node. Note that it was unfair to directly compare our protocol with Dumbo in terms of the average data usage per node per instance because the votes of a vote height counted in our protocol were also for other instances, and therefore, also counted in the other instances. Consequently, our data usage should be considered much smaller than Dumbo2 and Dumbo-NG.
@@ 7.3 Short Block Intervals
We evaluated the effects of shorter block intervals—1 and 0.5 seconds—on node synchronization (see Tbl. 3), using the same experimental setup as given in previous experiments. The findings indicated that it took 2 and 7 rounds to reach consensus under good-case latency, whereas under bad-case latency, it took 21 and 41 rounds, respectively. Here, the extended voting period for bad-case latency primarily served to wait for the slowest honest node’s vote. Notably, while nodes aligned to the same branch within a few rounds, reaching a consensus required receiving the votes from the slowest honest nodes.
Table 3: Experiment Results for shorter block interval (100 epochs with 1000 transactions emitted every 5 secs).
Latency Good-case Good-case Bad-case Bad-case
Block Interval (second) 1 0.5 1 0.5
Avg. Transaction Inclusion Rate 1 1 1 1
Rounds of voting 2 7 21 41
Avg. Block Acceptance Time (second) 2.51 3.88 21.82 20.75
Avg. Data Usage per Node per instance (MB) 0.87 2.71 8.36 17.34
@@
We did not test the case where an adversary constantly controlled entire network delays to such an extent that it could precisely and constantly decide which $N-f$ votes from the last height would be received first by each node. Even when using the leader-collecting system, where the leader could decide which votes to send, it would require the honest nodes always to recognize the adversary node as the leader node to perform this attack, which was impossible. The amendment to overcome this improbable attack was outlined in Sec. 4.6.2. We believed this attack abused the asynchronous network assumption and was not realistic, especially in the decentralized or at least federated networks.
8 Multiple Concurrent Leader Nodes
@@
The transaction inclusion blockages may occur if the volume of uncommitted transactions exceeds the capacity of the system. For example, when $\Delta BI=5$ and new transactions of a block size are emitted to each node also in 5 seconds, as set up in Sec. 7, the transactions are delayed in being included in the blockchain due to a faulty leader wasting 5 seconds of its leader time. As such, only around $2/3$ of transactions are included in the blockchain because only around $2/3$ of the leaders are honest.
@@
9 Conclusion
Our robust validated strong BFT consensus protocol is specially designed for leader-and-vote-based blockchains, particularly suited for asynchronous networks. It surpasses existing asynchronous blockchain solutions in terms of operational simplicity and consensus efficiency. With adversarial tolerance comparable to that of binary byzantine agreement, it ensures the seamless progression of a blockchain without requiring immediate consistency among nodes. By addressing key challenges in achieving termination and agreement properties within asynchronous environments, our protocol ensures liveness in consensus and significantly reduces message complexity associated with asynchronous BFT algorithms.
Appendix A Proofs
[…]
Appendix B Rounds Until Convergence
We present straightforward simulations demonstrating nodes arriving at a common branch. As detailed in the paper, nodes must contribute their votes to the branch with the highest support as suggested by the vote count (VC) of the votes, otherwise the votes are non-compliant votes and are not counted. We did simple simulations using a Python code provided in Code Listing 1 to test the average number of voting rounds (vote heights) used until all nodes converge on the same branch in practice. In each run of the simulation, each node votes for a random block among the $m$ blocks at the first round of voting. To simulate the random network delays in asynchronous network, each node will receive a random $N-f$ compliant set of votes. They then vote for the most supported one determined through a connected vote graph constructed from this random set of $N-f$ compliant votes. This simulation faithfully represents the normal scenarios where the adversary lacks control over the network delays. The adversary must also vote in the same way, as otherwise, their votes are not compliant votes. There is no need to consider other adversarial behaviours like double-voting, which is only a concern in the consensus-reestablishing process, not the process of reaching the same branch. Fig. 5 shows the simulation results for a $N=100$, $f=33$ system with varying initial numbers of blocks ($m$). Each simulation with an initial $m$ is repeated 100 times. Fig. 5 suggests that even when all nodes generated a block (initial $m=100$), it still took no more than five vote heights for all nodes to reach the same branch.
@@ Appendix B Rounds Until Convergence
Appendix C Discussion
Achieving an asynchronous SMR with minimal message and communication overhead is crucial, especially considering the inherently asynchronous nature of the Internet. Our protocol’s ability to match the message complexity of a binary partially synchronous leader-vote-based SMR (at the level of $O(N^{2})$ or $O(N)$ if the leader is responsible for transmitting votes), alongside its resilience in maintaining stable periodic state progression even in the presence of slow nodes and within the most relaxed communication model, represents a significant advancement in distributed systems. This accomplishment introduces a validated strong BFT consensus mechanism to the community, capable of aligning the complexity of the asynchronous state machine replication problem with that of the partially synchronous version. It not only enables the blockchain to seamlessly integrate the asynchronous BFT protocol, renowned for its efficiency without excessive data requirements or message complexity, but also strengthens the system against potential attacks associated with synchronous networks (e.g., eclipse attack [17], routing attack [3]), particularly those reliant on time-bound assumptions.

### v2

@@ 4.2.2 Comparison with other asynchronous blockchains
4.3 Optimal Two-Round Voting Process
HotStuff-2 requires three continuous block intervals to reach a consensus. During these intervals, all honest nodes are aligned, and messages are successfully exchanged before each interval concludes. Similarly, our protocol also necessitates three block intervals: one for proposing the block and two for voting. In the initial block interval (consensus height $k^{\prime}$), block $B_{k^{\prime}}$ is proposed and received by all before the interval ends. In the subsequent interval, all honest nodes cast their votes for $B_{k^{\prime}}$, and these votes are received by all before the interval’s close. By the third interval (vote height $k^{\prime}+2$), all honest nodes can verify through their vote counts that $B_{k^{\prime}}$ has garnered enough support during vote height $k^{\prime}+1$, thereby satisfying Eq. 2 and achieving consensus. Thus, our protocol aligns with the optimal voting rounds of HotStuff-2 and efficiently supports $O(N)$ view changes when the network is synchronous.
@@ 5 Implementation
We did not experiment with DAG-based protocols. They exhibited similar complexity to ACS-based protocols. Comparisons with DAG-based protocols are given in Appx. A.3.
@@
At the start of each 5 seconds, 1000 identical transactions (each sized 256 bytes) in total appeared in each node’s local database. This was regardless of the progress of the blockchains. The leader node grouped the first 1000 transactions not written to the blockchain to form a block. For nodes in Dumbo, they broadcasted the uncommitted transactions, which together sized $2.05\times 10^{6}$ bits. Further justifications of the experiment metrics are given in Appx. C.4.
@@ 5.1.1 Methodology
Regarding voting, adversarial nodes in our algorithm were constrained to make compliant votes. If they voted, they expedited consensus. We, therefore, assumed that adversarial nodes in our protocol and Hotstuff would attempt to create forks when selected as a leader, and otherwise kept silent, akin to the behavior of silent Byzantine nodes in Dumbo.
@@ 5.1.2 Results
Our protocol used a similar average block acceptance time compared to HotStuff-2, while HotStuff-2 required smaller data; both approaches largely outperformed Dumbo2 and Dumbo-NG due to reduced complexity, as expected. Tbl. 2 reported the experiment results.
@@ 5.2.2 Results
As expected, our protocol largely outperformed Dumbo2 and Dumbo-NG due to reduced complexity. Tbl. 3 presented the experimental results. The transaction inclusion rate for our protocol was similar in both good and bad-case latency because the only factors that affected it in our experiment were the rate of adversarial leaders and the pre-set block interval. The average data usage per node per instance increased because rounds of voting occurred while awaiting the votes from the slowest honest node. Note that it was unfair to directly compare our protocol with Dumbo in terms of the average data usage per node per instance because the votes of a vote height counted in our protocol were also for other instances, and therefore, also counted in the other instances. Consequently, our data usage should be considered much smaller than Dumbo2 and Dumbo-NG.
@@ 5.3 Short Block Intervals
We evaluated the effects of shorter block intervals—1 and 0.5 seconds—on node synchronization (see Fig. 5), using the same experimental setup as given in previous experiments. The findings indicated that it took 2 and 7 rounds to reach consensus under good-case latency, whereas under bad-case latency, it took 21 and 41 rounds, respectively. Here, the extended voting period for bad-case latency primarily served to wait for the slowest honest node’s vote. Notably, while nodes aligned to the same branch within a few rounds, reaching a consensus required receiving the votes from the slowest honest nodes.
Figure 5: Experiment Results for shorter block interval (100 epochs with 1000 transactions emitted every 5 secs).
@@
We did not test the case where the adversary constantly controlled entire network delays to such an extent that it could precisely and constantly decide which $N-f$ votes from the last height would be received first by each node. Even when using the leader-collecting system, where the leader could decide which votes to send, it would require the honest nodes always to recognize the Byzantine node as the leader node to perform this attack, which was impossible. The amendment to overcome this improbable attack was outlined in Sec. 3.6.2. We believed this attack abused the asynchronous network assumption and was not realistic, especially in the decentralized or at least federated networks.
6 Conclusion
Our robust validated strong BFT consensus protocol is specially designed for leader-and-vote-based blockchains, particularly suited for asynchronous networks. It surpasses existing asynchronous blockchain solutions in terms of operational simplicity and consensus efficiency. With adversarial tolerance comparable to that of binary byzantine agreement algorithms, it ensures the seamless progression of a blockchain without requiring immediate consistency among nodes. By addressing challenges in achieving termination and agreement properties within asynchronous environments, our protocol ensures liveness in consensus and significantly reduces message complexity associated with asynchronous BFT algorithms.
Appendix A Related Works
A.1 Strong BFT Consensus Protocols
Safety. A Byzantine node may double vote or send its votes only to a subset of honest nodes. However, this cannot trick some nodes into believing that an input value has been advanced to a new round when it actually did not. This is because less than $N-f$ nodes will confirm that the input value has received more than $f$ votes. There could be different sets of $N-f$ reports, resulting in some nodes considering that an input value was not confirmed, even though it was confirmed. But, these nodes will work this out eventually during the third phase as there must be more than $f$ nodes who have accepted this input value into the next round. So, after each voting round, the honest nodes reach the same set of remaining input values.
A.1.1 Challenges in guaranteeing liveness of consensus derivation with the leader-based design
The strong BFT consensus protocol in Appx. A.1 requires a known set of $m$ input values and tolerates only a small $f$. It does not consider the origin of the input values. If $m$ leaders propose them, then consensus liveness in asynchronous settings is not guaranteed unless $m\geq N-f$. Also, when $m>2$ and $f\leq\lfloor(N-1)/3\rfloor$, it is possible that no input value receives more than $f$ votes, making it impossible to discard a block that doesn’t receive $f+1$ votes.
In contrast, our protocol does not require a known $m$ but ensures that, via multiple voting rounds, honest nodes will converge to the same value.
A.2 ACS-based Protocols
A.3 DAG-based Protocols
Direct Acyclic Graph (DAG)-based BFT protocols [19, 10, 26] use DAGs to transmit transactions. Each node sends a block linked to $N-f$ blocks from the preceding height. A leader block is periodically selected to order and finalize the linked blocks. To avoid equivocation, they must reliably broadcast each block, driving the overall complexity to $O(N^{3})$. A promising recent preprint [18] aims to reach consensus while allowing equivocation in asynchronous networks during the process, thereby reducing the overall complexity.
DAG-based protocols are particularly useful for avoiding delays caused by slow or faulty leaders. Instead of selecting a leader first and waiting for the leader node to propose a block of transactions for each block height, every node proposes a batch of transactions it knows, with the leader node selected later and the batch of the leader node be finalized as the consensus. This approach prevents delays caused by a pre-determined leader node but comes with the trade-off of increased data usage. Since each block is linked to $N-f$ blocks from the last round, each block can also be viewed as a vote for the previous rounds.
Our protocol is very similar to DAG-based protocols if all $N$ nodes propose a block at each block height. In this setting, our protocol could also be used to avoid the delays caused by faulty or slow leader nodes. However, there are significant differences even in this setting:
  1. 1.
    The method of achieving consensus is entirely different. DAG-based protocols still require a deterministic leader, albeit with the leader election postponed until after nodes broadcast their blocks. To ensure a deterministic leader, they must determine the exact blocks broadcast to avoid equivocation. Optimizations so far only seek to patch equivocation issues without involving Byzantine reliable broadcasting. In contrast, our approach does not require a deterministic leader at all. Instead, we leverage the vote graph to gradually force all nodes to align with the same branch.
  2. 2.
    Since votes and blocks are separated in our approach, we require all nodes to wait until they receive $N-f$ votes instead of $N-f$ blocks. The size of blocks is significantly larger than the size of votes. Thus, our protocol should be much faster than DAG-based protocols and demand less data.
  3. 3.
    Nodes in our protocol do not need to download all $N-f$ blocks per round for the protocol to function. Instead, each node only needs to download the chain of blocks leading up to the block it votes for. As such, unless it needs to switch to another branch, each node downloads only one block per height in normal cases, regardless of how many other blocks are present. In Appx. D, we show that even when all nodes propose a block, nodes converge in only few rounds. Thus, nodes only switch branches few times before a block is consensually accepted. This supports our claim of significantly reduced data usage and significantly less message complexity compared to DAG-based protocols.
Thus, despite also using graphs, our approach is much more similar to classical leader-based methods such as PBFT and HotStuff where nodes only download a block that they vote for, making our protocol distinct from current DAG-based approaches while requiring much less data.
Appendix B Proofs
[…]
Appendix C Further Analysis and Experiment
C.1 Remarks on the FLP Result
The Fischer-Lynch-Paterson (FLP) result [13] demonstrates that achieving agreement in an asynchronous message-passing system is impossible if even a single crash failure is allowed, unless the basic model is augmented with additional mechanisms, such as randomization or failure detectors.
Our protocol addresses this limitation through two key mechanisms:
  1. 1.
    Randomness from network delays: The inherent unpredictability of network delays allows nodes to construct different vote graphs randomly, influencing their voting decisions in subsequent rounds.
  2. 2.
    Common coin flip: In cases where the randomness of network delays is insufficient or absent, a common coin flip is employed as a source of randomization.
These features ensure that our protocol incorporates sufficient randomization to circumvent the constraints imposed by the FLP result. Consequently, the correctness of our protocol is not affected by the FLP impossibility.
C.2 Remarks on Consensus Chain Propagation
In Sec. 3.3, we assumed that nodes have access to a local clock. However, this assumption does not imply that the system is partially synchronous or synchronous in terms of communication. The local clock is neither established through communication among nodes nor required to align perfectly. The sole purpose of this assumption is to provide a trigger for certain nodes to act as leader nodes, proposing blocks periodically.
Since the uniqueness of the leader node is not required, this mechanism can be achieved through other means of randomness. For example, a node could monitor its own CPU speed, which is ostensibly unpredictable. When the CPU speed coincidentally matches a pre-set random number, the node proposes a new block. This approach also helps ensure liveness by facilitating the extension of the consensus chain.
C.3 Remarks on Vote Compliance Checking
It is well known that a perfect Byzantine failure detector does not exist [6, 12]. The core issue lies in the fact that, without guaranteeing that all honest nodes observe the same information, they cannot make consistent judgments. Indeed, in asynchronous networks, there is no assurance that honest nodes will observe the same information, not even eventually, as nodes may send different messages to different recipients.
However, our protocol focuses on detecting non-compliant votes rather than Byzantine nodes. Specifically, nodes are required to justify their vote for a particular block by demonstrating that this block belongs to the most supported branch according to the vote graph derived from the vote. When each node receives a vote, they can construct identical vote graphs using this vote, making it straightforward for all honest nodes to reach the same judgment based on the information in the vote graph.
Furthermore, as each vote graph is constructed using the $N-f$ vote graphs from the previous height—many of which originate from honest nodes—the adversary’s voting decisions are progressively constrained round by round.
Thus, our protocol does not aim to detect Byzantine nodes directly. Instead, it systematically reduces the valid choices for nodes in each round through vote compliance checking.
C.4 Remarks on Experimental Metrics
Our experimental metrics differ from those used in existing asynchronous BFT consensus protocols, which measure transactions accepted per second (TPS) with varying batch/block sizes. Directly relating the batch size of an ACS-based protocol to TPS suggests a hidden assumption that transactions among all batches are distinct. This is unrealistic in a real-world blockchain where users may submit their transactions to multiple nodes in an attempt to have their transactions included (because they do not know if a particular node is honest) in the blockchain. Our experimental setup, thus, considers a realistic scenario in which 1,000 transactions are emitted to all nodes every 5 seconds, causing large overlap in the batches sent by each node in ACS-based protocols.
In our protocol, block proposals are entirely decoupled from the voting process, allowing new blocks to be broadcast at regular, fixed intervals, regardless of whether the previous block has been accepted. Variations in block intervals can affect the number of branches in the network. Because leader nodes extend the consensus chain known to themselves, different consensus chains with a common prefix can develop into branches. If new blocks are proposed at a very fast rate, it will result in a great number of branches before nodes convergence into the same one. This scenario provides an interesting experimental setup for evaluating performance across different block intervals and other factors that might influence branch formation.
The batch/block size significantly influences latency (the time taken to accept transactions) in ACS-based or DAG-based protocols. This distinction arises because, in those protocols, every node must broadcast its set of transactions and each node must receive at least $N-f$ batches of transactions per instance; thus, small changes in batch size can drastically impact overall data usage, affecting both the number of transactions per batch and the latency.
In contrast, in our protocol, only the leader node broadcasts a block under normal circumstances. Ideally, a node only downloads $N$ votes and a block per height. A node is required to download other blocks only when it switches to vote in different branches. Even if adversaries broadcast more blocks and create additional branches, each node only synchronizes the chain of blocks leading up to the block it votes for when switching branches. Consequently, the batch/block size has rather minimal impact on experimental outcomes. Instead, the frequency with which nodes switch to vote for different branches affects performance.
Thus, experimenting with different batch/block sizes becomes uninteresting and not very insightful in terms of TPS variations. With that said, in our experiment, $TPS=\frac{\text{Transactions in a block}}{\text{Avg. Block Acceptance Time}}$.
C.5 Multiple Concurrent Leader Nodes
@@
The transaction inclusion blockages may occur if the volume of uncommitted transactions exceeds the capacity of the system. For example, when $\Delta BI=5$ and new transactions of a block size are emitted to each node also in 5 seconds, as set up in Sec. 5, the transactions are delayed in being included in the blockchain due to a faulty leader wasting 5 seconds of its leader time. As such, only $2/3$ of transactions are included in the blockchain because only $2/3$ of the leaders are honest.
@@
Appendix D Rounds Until Convergence
We present straightforward simulations demonstrating nodes arriving at a common branch. As detailed in the paper, nodes must contribute their votes to the branch with the highest support as suggested by the vote count ($\operatorname{VC}$) of the votes, otherwise the votes are non-compliant votes and are not counted. We did simple simulations using a Python code provided in Code Listing 1 to test the average number of voting rounds (vote heights) used until all nodes converge on the same branch in practice. In each run of the simulation, each node votes for a random block among the $m$ blocks at the first round of voting. To simulate the random network delays in asynchronous network, each node will receive a random $N-f$ compliant set of votes. They then vote for the most supported one determined through a connected vote graph constructed from this random set of $N-f$ compliant votes. This simulation faithfully represents the normal scenarios where the adversary lacks control over the network delays. The Byzantine nodes must also vote in the same way, as otherwise, their votes are not compliant votes. Fig. 6 shows the simulation results for a $N=100$, $f=33$ system with varying initial numbers of blocks ($m$). Each simulation with an initial $m$ is repeated 100 times. Fig. 6 suggests that even when all nodes generated a block (initial $m=100$), it still took no more than five vote heights for all nodes to reach the same branch.
@@ Appendix D Rounds Until Convergence
Figure 8: The visualization of Fig. 7. $(X,Y)$ refers to the vote for block ID $Y$ at height $X$. Convergence reached at height 3.
@@
Figure 10: The visualization of Fig. 9. $(X,Y)$ refers to the vote for block ID $Y$ at height $X$. Convergence reached at height 4
@@
Figure 12: The visualization of Fig. 11. $(X,Y)$ refers to the vote for block ID $Y$ at height $X$. Convergence reached at height 4
@@
Figure 14: The visualization of Fig. 13. $(X,Y)$ refers to the vote for block ID $Y$ at height $X$. Convergence reached at height 3
Appendix E Using the Validated Strong BFT Consensus Visualization Tool
To help understand our Validated Strong BFT Consensus protocol, we provide a toy experiment as well as its visualization tool for illustration, which visualizes the protocol in action, demonstrating how BFT consensus is achieved in our protocol in an asynchronous environment.
The interactive version of our tool can be accessed via the following anonymous link: https://waterandcola.github.io/validated-strong-consensus/index.html
Steps to Try the Tool:
1. Access the Webpage: Navigate to the provided URL to open the interactive tool in your browser. No local setup or additional dependencies are needed.
2. Replay the Consensus Process: The example provided simulates a consensus network of 7 nodes ($N=7$) with a fault tolerance of up to 2 faulty nodes ($f=2$). Each node proposes blocks and casts votes in asynchronous rounds to reach consensus. In this process, all 7 nodes behave normally. This is because if the adversary cast non-compliant votes, their votes are not processed, resulting in 5 active honest nodes and two silent Byzantine nodes. Since each node only votes after receiving $N-f$ votes from the previous height, having only 5 active honest nodes leads to very fast branch alignment (since each one will have the same view after one voting round), making it less interesting for illustration.
3. View Logs and Replay Consensus:
4. Visualize the Consensus Graph:
  • The tool generates a dynamic graph where blocks (squares) and votes (circles) are displayed.
  • Green blocks represent blocks that have reached consensus after accumulating the required number of votes.
  • You can hover over votes to see additional information about the vote, voter graph, vote counter and the blocks they support, as shown in Fig. 15.
    Figure 15: An explanation for visualization tool (https://waterandcola.github.io/validated-strong-consensus/index.html).